A project has an initial cost of $200,000 and uniform annual benefits of $35,000. At the end of its 8-year useful life, its salvage value is $50,000. At a 10% interest rate, the net present worth of the project is approximately what amount?

       
$30,250

       
$130,000

       
$10,050

       
$34,825
Answer #1

The net present worth of the project is:-

= -200000+35000*(P/A,10%,8)+50000*(P/F,10%,8)

= -200000+35000*5.334926+50000*0.466507

= -200000+186722.4+23325.37

= 10048

Option(C)